# Finite-Sample Non-Parametric Bounds with an Application to the Causal Effect of Workforce Gender Diversity on Firm Performance

### v1

@@ 3 Estimation and Identification
The estimator of $\delta$ can be written as a single weighted sum:
$\hat{\delta}=\sum_{i=1}^{n^{j}}\sum_{j=1}^{K}\sum_{t=1}^{T}Y_{ijt}\,w_{ijt},\qquad w_{ijt}=\frac{Z_{ijt}}{N_{1}}\;-\;\frac{1-Z_{ijt}}{N_{0}},$
@@
$N_{k}=\sum_{i,j,t}\mathbbm{1}\{Z_{ijt}=k\},\quad k\in\{0,1\},\quad N=N_{0}+N_{1}.$
@@
To estimate $\Re$, recall from (14) that $\Re$ involves the four quantities $\mathbb{E}[Y^{(k)}\mid Z=k]$ and $\Pr(Z=k)$, $k=0,1$, plus the endpoints $\{L^{(k)},U^{(k)}\}$. We estimate them by
$\hat{\delta}_{k}=\frac{1}{N_{k}}\sum_{i,j,t}Y_{ijt}\,\mathbbm{1}\{Z_{ijt}=k\},\qquad\hat{p}_{k}=\frac{N_{k}}{N},$
@@
$\hat{F}^{(k)}(y)=\frac{1}{N_{k}}\sum_{i,j,t}\mathbbm{1}\{Y_{ijt}\leq y,\;Z_{ijt}=k\},\qquad\hat{Q}_{Y}^{(k)}(p)=\inf\{y:\hat{F}^{(k)}(y)\geq p\},$
@@ 4 Inference
For $u=m_{0},\cdots,m_{1}$, obtaining the $(1-\alpha_{u})\%$ confidence intervals for the naïve estimator $\hat{\delta}$ is rather straightforward, since $\hat{\delta}$ is a linear statistic and from (19) it follows that:
$\sigma^{2}(\hat{\delta})=\mathbb{E}\left[(Y_{ijt}w_{ijt}-\delta)^{2}\right],$ (20)
@@
$\hat{\sigma}^{2}(\hat{\delta})=\frac{1}{N-1}\sum_{i,j,t}\left(Y_{ijt}w_{ijt}-\hat{\delta}\right)^{2}.$ (21)
A $(1-\alpha_{u})\%$ Wald‐type interval is then
$\hat{\delta}\pm\Phi^{-1}(1-\alpha_{u}/2)\sqrt{\hat{\sigma}^{2}(\hat{\delta})},\quad\text{for}\quad u=m_{0},\cdots,m_{1},$
where $\Phi^{-1}(.)$ is the inverse of CDF of the standard normal distribution.
Obtaining confidence interval for the nonparametric bounds is less straightforward, since the upper/lower estimators are nonlinear. As such we have the following proposition which is predicated on Bonferroni-adjusted delta-method.
@@ A $(1-\alpha_{u})\%$ Wald‐type interval is then
$\nabla\mathcal{L}(\theta)=\left(p_{1},-p_{0},\delta_{1}-U_{0},L_{1}-\delta_{0}\right)^{\top}$ (23)
$\nabla\mathcal{U}(\theta)=\left(p_{1},-p_{0},\delta_{1}-L_{0},U_{1}-\delta_{0}\right)^{\top}$ (24)
and the covariance of the estimators $\hat{\theta}$ is given explicitly by:
@@ 4.1 Concentration-Driven Confidence Bands for Average Treatment Effects
A major shortcoming of the nonparametric bounds proposed by Manski, (1990, 2003) and introduced in Section 2 is the strong assumption that the latent conditional expectations in Eq. (9) lie inside a known bounded interval
$\left[L^{(k)},\,U^{(k)}\right],\quad k\in\{0,1\}.$
In practice, these expectations may be unbounded. To address this, we reformulate the problem probabilistically and study
$\Pr\left(\forall u\in\mathcal{M},\Re_{u}\in H_{\alpha_{u}}[\Re_{u}]\right).$
We first analyze the stylized case in which the observations $Y_{ijt}$ are independent for all indices $i,j,t$. Although independence is useful for pure cross‑sectional snapshots, where firms may be assumed to be independent units—for example, analyzing one quarter across many sectors—it is clearly unrealistic in panel settings, where serial correlation may be present. Consequently, we extend our results to the more realistic scenario in which each firm‑level time series $(Y_{ij1},\cdots,Y_{ijT})$ exhibits temporal dependence. In essence, we allow temporal dependence within firm, but assume firms are independent cross-sectionally.
Simplifying the notations in section 2, identification follows Manski, (1990, 2003):
$\Re\in\left[\delta_{1}p_{1}+L^{(1)}p_{0}-U^{(0)}p_{1}-\delta_{0}p_{0},\;\delta_{1}p_{1}+U^{(1)}p_{0}-L^{(0)}p_{1}-\delta_{0}p_{0}\right],$ (26)
where $\delta_{k}:=\mathbb{E}[Y^{(k)}_{ijt}\mid Z_{ijt}=k]$ and $p_{k}:=\Pr(Z_{ijt}=k)$. Replacing the latent terms $(\delta_{10},\delta_{01})$ (where $\delta_{kk^{\prime}}:=\mathbb{E}[Y_{ijt}^{(k)}\mid Z_{ijt}=k^{\prime}]$) by support limits $(L^{(k)},U^{(k)})$ yields (26). In the first setting, we assume that the data exhibits no cross-sectional or serial dependence.
@@
Proposition 2 (Finite–sample coverage under i.i.d. sampling).
Let $0<\alpha_{u}<1$ for $u=m_{0},\cdots,m_{1}$ and denote by $N_{k}=\sum_{i}\mathbbm{1}\{Z_{i}=k\}$ the sample size in treatment group $k\in\{0,1\}$ and by
@@
$\varepsilon_{k}:=\sqrt{\frac{\log\left(12/\alpha_{u}\right)}{2N_{k}}},\quad L_{\alpha_{u}}^{(k)}:=Y^{(k)}_{(1)}-\varepsilon_{k},\quad U_{\alpha_{u}}^{(k)}:=Y^{(k)}_{(N_{k})}+\varepsilon_{k},\qquad k=0,1.$
and define the two thresholds
@@
$\begin{split}H_{\alpha_{u}}[\Re_{u}]=\left[\hat{\mu}_{1}^{-}\hat{p}_{1}^{-}+L_{\alpha_{u}}^{(1)}\hat{p}_{0}^{-}-U_{\alpha_{u}}^{(0)}\hat{p}_{1}^{+}-\hat{\mu}_{0}^{+}\hat{p}_{0}^{+},\;\hat{\mu}_{1}^{+}\hat{p}_{1}^{+}+U_{\alpha_{u}}^{(1)}\hat{p}_{0}^{+}-L_{\alpha_{u}}^{(0)}\hat{p}_{1}^{-}-\hat{\mu}_{0}^{-}\hat{p}_{0}^{-}\right],\end{split}$ (27)
where $\hat{\mu}_{k}^{\pm}=\hat{\mu}_{k}\pm t_{\mu,k}$ and $\hat{p}_{k}^{\pm}=\hat{p}_{k}\pm t_{p,k}$ Then, under assumption 1
$\Pr\left(\forall u\in\mathcal{M},\Re_{u}\in H_{\alpha_{u}}[\Re_{u}]\right)\geq 1-\alpha,$ (28)
Proposition 2 states that the data–driven set $H_{\alpha_{u}}[\Re_{u}]$ in (29) is a $100(1-\alpha_{u})\%$-level confidence region for the average treatment effect $\Re_{u}$ under nothing more than i.i.d. sampling. Because $\mathbb{E}[Y_{ijt}^{(1)}\mid Z_{ijt}=0]$ and $\mathbb{E}[Y_{ijt}^{(0)}\mid Z_{ijt}=1]$ are latent, point identification is impossible without additional assumptions; the proposition nevertheless guarantees that the random interval constructed from the empirical means, treatment proportions, and slightly “padded” sample extrema will cover the true $\Re_{u}$ in at least $100(1-\alpha_{u})\%$ of repeated samples. Practically, one computes $H_{\alpha}[\Re]$ by (i) splitting the sample into treated and untreated subsamples, (ii) forming the subsample–specific means $\mu_{k}$ and proportions $\hat{p}_{k}$, (iii) widening the minimal and maximal observed outcomes by the DKWM envelope $\varepsilon_{k}=\sqrt{\log(12/\alpha_{u})/2N_{k}}$, and (iv) plugging these objects into (29). The resulting band can be used exactly like an ordinary confidence interval: the null hypothesis $H_{0}:\Re_{u}=r_{0}$ is rejected at level $\alpha_{u}$, whenever $r_{0}\notin H_{\alpha_{u}}[\Re_{u}]$.
If substantive knowledge implies that the latent outcomes are truncated on one or both tails—for instance, Tobin’s $Q$ is bounded below by 0—the extreme-value inputs in Manski’s bounds can be replaced by the true population limits. Let $\lambda$ (lower) and $\Lambda$ (upper) denote any such known bounds. When both limits are known one sets $a=\lambda$ and $b=\Lambda$ in the plug-in formulas; the resulting $100(1-\alpha)\%$ simultaneous band coincides with Proposition 1 and requires no DKW padding. When only one tail is known—say $Y\geq\lambda$, but the upper support is unknown—we fix the lower extreme at $\lambda$ while retaining the sample maximum, the DKW envelope on the upper side. The next Corollary shows that this hybrid construction preserves the nominal family-wise coverage probability even when the first significant threshold is data-selected.
Corollary 1 (Finite–sample coverage under i.i.d. sampling and truncated distribution).
Let $0<\alpha_{u}<1$ for $u=m_{0},\cdots,m_{1}$ and denote by $N_{k}=\sum_{i}\mathbbm{1}\{Z_{i}=k\}$ the sample size in treatment group $k\in\{0,1\}$ and by
$\lambda\leq Y^{(k)}_{(1)}\leq\cdots\leq Y^{(k)}_{(N_{k})}$
the order statistics of the observed outcomes in that group. Set
$\varepsilon_{k}:=\sqrt{\frac{\log\left(6/\alpha_{u}\right)}{2N_{k}}},\quad L^{(k)}:=\lambda,\quad U_{\alpha_{u}}^{(k)}:=Y^{(k)}_{(N_{k})}+\varepsilon_{k},$
for $k=0,1$, and define the two thresholds
$\begin{split}H_{\alpha_{u}}[\Re_{u}]=\left[\hat{\mu}_{1}^{-}\hat{p}_{1}^{-}+L^{(1)}\hat{p}_{0}^{-}-U_{\alpha_{u}}^{(0)}\hat{p}_{1}^{+}-\hat{\mu}_{0}^{+}\hat{p}_{0}^{+},\;\hat{\mu}_{1}^{+}\hat{p}_{1}^{+}+U_{\alpha_{u}}^{(1)}\hat{p}_{0}^{+}-L^{(0)}\hat{p}_{1}^{-}-\hat{\mu}_{0}^{-}\hat{p}_{0}^{-}\right],\end{split}$ (29)
where $\hat{\mu}_{k}^{\pm}=\hat{\mu}_{k}\pm t_{\mu,k}$ and $\hat{p}_{k}^{\pm}=\hat{p}_{k}\pm t_{p,k}$ Then, under assumption 1
$\Pr\left(\forall u\in\mathcal{M},\Re_{u}\in H_{\alpha_{u}}[\Re_{u}]\right)\geq 1-\alpha,$ (30)
@@
Proposition 3 (Finite–sample coverage under $\upalpha$–mixing sampling).
Let $0<\alpha_{u}<1$ for $u=m_{0},\cdots,m_{1}$ and let $(Y_{ijt},Z_{jt})_{i,j,t}$ be a strictly stationary sequence with $Z_{i}\in\{0,1\}$, $Y_{i}\in\mathbb{R}$, and strong–mixing coefficients $\upalpha(r)$ satisfying
@@
$N_{k}=\sum_{i=1}^{n}\mathbbm{1}\{Z_{i}=k\},\quad\hat{p}_{k}=\frac{N_{k}}{N_{0}+N_{1}},\quad\hat{\mu}_{k}=\frac{1}{N_{k}}\sum_{Z_{i}=k}Y_{i},\quad k=0,1.$
@@
$t_{p,k}=(1+4C_{\upalpha})\sqrt{\frac{2\log(12/\alpha_{u})}{N_{k}}},\qquad t_{\mu,k}=\max\left\{t_{k}^{(1)},t_{k}^{(2)},t_{k}^{(3)}\right\},$
@@
$\varepsilon_{k}=t_{p,k},\quad L^{(k)}_{\alpha_{u}}=Y^{(k)}_{(1)}-\varepsilon_{k},\quad U^{(k)}_{\alpha_{u}}=Y^{(k)}_{(N_{k})}+\varepsilon_{k},\quad k=0,1.$
@@
$H_{\alpha_{u}}[\Re_{u}]=\left[\hat{\mu}_{1}^{-}\hat{p}_{1}^{-}+L^{(1)}_{\alpha}\hat{p}_{0}^{-}-U^{(0)}_{\alpha}\hat{p}_{1}^{+}-\hat{\mu}_{0}^{+}\hat{p}_{0}^{+},\;\hat{\mu}_{1}^{+}\hat{p}_{1}^{+}+U^{(1)}_{\alpha}\hat{p}_{0}^{+}-L^{(0)}_{\alpha}\hat{p}_{1}^{-}-\hat{\mu}_{0}^{-}\hat{p}_{0}^{-}\right]$
where $\hat{\mu}_{k}^{\pm}=\hat{\mu}_{k}\pm t_{\mu,k}$ and $\hat{p}_{k}^{\pm}=\hat{p}_{k}\pm t_{p,k}$ Then, under assumption 2
$\Pr\left(\forall u\in\mathcal{M},\Re_{u}\in H_{\alpha_{u}}[\Re_{u}]\right)\geq 1-\alpha.$
Similar to Corollary 1, the extension of Proposition 3 to the case of one-tail truncated latent conditional expectations simply requires modifying the paddings to the bounds. Specifically, for the case of lower tail truncation, replace
$L_{\alpha_{u}}^{(k)}:=Y_{(1)}^{(k)}-\varepsilon_{k}\quad\text{and}\quad U^{(k)}_{\alpha_{u}}:=Y_{(N_{k})}^{(k)}+\varepsilon_{k}$
with
$L^{(k)}:=\lambda,\quad\text{and}\quad U^{(k)}_{\alpha_{u}}:=Y_{(N_{k})}^{(k)}+\tilde{\varepsilon}_{k},$
where now
$\tilde{\varepsilon}_{k}:=(1+4C_{\upalpha})\sqrt{\frac{2\log\left(6/\alpha_{u}\right)}{N_{k}}}$
for $k=0,1$, where the observed order statistics satisfy $\lambda\leq Y_{(1)}^{(k)}\leq\cdots\leq Y_{(N_{k})}^{(k)}$.

### v2

@@ 3 Estimation and Identification
For a fixed $\tau$, the estimator of $\delta$ can be written as a single weighted sum:
$\hat{\delta}=\sum_{i=1}^{n^{j}}\sum_{j=1}^{K}\sum_{t=1}^{T}Y_{ijt}\,w_{ijt},\qquad w_{ijt}=\frac{Z_{ijt}(\tau)}{N_{1}}-\frac{1-Z_{ijt}(\tau)}{N_{0}},$
@@
$N_{k}=\sum_{i,j,t}\mathbbm{1}\{Z_{ijt}(\tau)=k\},\quad k\in\{0,1\},\quad N=N_{0}+N_{1}.$
@@
To estimate $\Re$, recall from (14) that $\Re$ involves the four quantities $\mathbb{E}[Y_{ijt}^{(k)}\mid Z_{ijt}(\tau)=k]$ and $\Pr(Z_{ijt}(\tau)=k)$, $k=0,1$, plus the endpoints $\{L^{(k)},U^{(k)}\}$. We estimate them by
$\hat{\delta}_{k}=\frac{1}{N_{k}}\sum_{i,j,t}Y_{ijt}\mathbbm{1}\{Z_{ijt}(\tau)=k\},\qquad\hat{p}_{k}=\frac{N_{k}}{N},$
@@
$\hat{F}^{(k)}(y)=\frac{1}{N_{k}}\sum_{i,j,t}\mathbbm{1}\{Y_{ijt}\leq y,\;Z_{ijt}(\tau)=k\},\qquad\hat{Q}_{Y}^{(k)}(p)=\inf\{y:\hat{F}^{(k)}(y)\geq p\},$
@@ 4 Inference
For $u=m_{0},\dots,m_{1}$, constructing $(1-\alpha_{u})$ confidence intervals for the naïve estimator $\hat{\delta}$ is straightforward because $\hat{\delta}$ is a difference of sample means. Recall from (21) that
$\sqrt{N}(\hat{\delta}-\delta)\xrightarrow{d}N(0,\sigma^{2}).$
@@
$\sigma^{2}(\hat{\delta})=\frac{\text{Var}(Y_{ijt}\mid Z_{ijt}(\tau_{u})=1)}{p_{1}}+\frac{\text{Var}(Y_{ijt}\mid Z_{ijt}(\tau_{u})=0)}{p_{0}}$ (22)
is the asymptotic variance parameter and $p_{k}=\Pr(Z_{ijt}(\tau_{u})=k)$. Hereafter, for notational simplicity and without loss of generality, we write $N_{k}$ instead of $N_{k}(\tau_{u})$ for each look $u$, keeping the dependence on $\tau_{u}$ implicit.
[…]
$\hat{\sigma}^{2}(\hat{\delta})=\frac{\hat{\sigma}_{1}^{2}}{\hat{p}_{1}}+\frac{\hat{\sigma}_{0}^{2}}{\hat{p}_{0}},$ (23)
where
$\hat{\sigma}_{k}^{2}=\frac{1}{N_{k}-1}\sum_{Z_{ijt}(\tau)=k}(Y_{ijt}-\bar{Y}_{k})^{2},\qquad\bar{Y}_{k}=\frac{1}{N_{k}}\sum_{Z_{ijt}(\tau)=k}Y_{ijt},$
and $\hat{p}_{k}=N_{k}/N$.
Since $\text{Var}(\hat{\delta})\approx\sigma^{2}(\hat{\delta})/N$, a $(1-\alpha_{u})$ Wald-type confidence interval for $\delta$ is therefore
$\hat{\delta}\pm\Phi^{-1}(1-\alpha_{u}/2)\,\sqrt{\frac{\hat{\sigma}^{2}(\hat{\delta})}{N}},\qquad u=m_{0},\dots,m_{1},$ (24)
where $\Phi^{-1}(\cdot)$ denotes the standard normal quantile function.
Obtaining confidence intervals for the nonparametric bounds is less straightforward, since the upper and lower bound estimators are nonlinear functions of the data. We therefore rely on a Bonferroni-adjusted delta method, as formalized in the following proposition.
@@
and the covariance matrix of the vector of estimators $\hat{\theta}$ is given explicitly by:
@@ 4.1 Concentration-Driven Confidence Bands for Average Treatment Effects
A major shortcoming of the nonparametric bounds proposed by manski1990nonparametric; manski2003partial and introduced in Section 2 is the strong assumption that the latent conditional expectations in Eq. (9) lie inside a known bounded interval
$L^{(k)},\;U^{(k)}],\qquad k\in\{0,1\}.$
In practice, these expectations may be unbounded. To address this, we reformulate the problem probabilistically and study the probability
$\Pr\left(\forall u\in\mathcal{M},\;\Re_{u}\in\mathcal{H}_{\alpha_{u}}[\Re_{u}]\right).$
We first analyse a stylized setting in which the observations $Y_{ijt}$ are independent across all indices $i,j,t$. Although independence may be reasonable for purely cross-sectional snapshots (for instance, a single quarter across many sectors) it is unrealistic in panel settings, where serial correlation is typically present. Accordingly, we later extend our results to allow for temporal dependence within firm, while maintaining cross-sectional independence across firms.
For each threshold value $\tau_{u}$, we define
$\delta_{k}:=\mathbb{E}\left[Y^{(k)}_{ijt}\mid Z_{ijt}(\tau_{u})=k\right],\qquad p_{k}:=\Pr\left(Z_{ijt}(\tau_{u})=k\right),\qquad k\in\{0,1\},$
with the dependence on $u$ left implicit to lighten notation. Following manski1990nonparametric; manski2003partial, the identified set for the tipping-point functional $\Re_{u}$ is then
$\Re_{u}\in\left[\delta_{1}p_{1}+L^{(1)}p_{0}-U^{(0)}p_{1}-\delta_{0}p_{0},\;\;\delta_{1}p_{1}+U^{(1)}p_{0}-L^{(0)}p_{1}-\delta_{0}p_{0}\right].$ (29)
The latent cross-terms $\delta_{10}:=\mathbb{E}[Y^{(1)}_{ijt}\mid Z_{ijt}(\tau_{u})=0]$ and $\delta_{01}:=\mathbb{E}[Y^{(0)}_{ijt}\mid Z_{ijt}(\tau_{u})=1]$ are replaced by their support bounds $(L^{(k)},\;U^{(k)})$ to obtain the interval in (29). In the first setting, we assume that the data exhibit neither cross-sectional nor serial dependence.
@@
Proposition 2 (Finite-sample coverage under i.i.d. sampling).
Let $0<\alpha_{u}<1$ for $u=m_{0},\cdots,m_{1}$ and denote by $N_{k}=\sum_{i}\mathbbm{1}\{Z_{i}(\tau_{u})=k\}$ the sample size in treatment group $k\in\{0,1\}$ and by
@@
$\varepsilon_{k}:=\sqrt{\frac{\log\left(12/\alpha_{u}\right)}{2N_{k}}},$
choose any $r_{k}\in\left(0,\frac{1}{2}-\varepsilon_{k}\right]$, and define
$L_{\alpha_{u}}^{(k)}:=Y^{(k)}_{\left(\lceil(r_{k}+\varepsilon_{k})N_{k}\rceil\right)},\quad U_{\alpha_{u}}^{(k)}:=Y^{(k)}_{\left(\lceil(1-r_{k}-\varepsilon_{k})N_{k}\rceil\right)},\quad k=0,1.$
(In particular, if $\varepsilon_{k}\leq\frac{1}{4}$, one may take $r_{k}=\varepsilon_{k}$, which yields $L_{\alpha_{u}}^{(k)}=Y^{(k)}_{(\lceil 2\varepsilon_{k}N_{k}\rceil)}$ and $U_{\alpha_{u}}^{(k)}=Y^{(k)}_{(\lceil(1-2\varepsilon_{k})N_{k}\rceil)}$.)
Define the two thresholds
Let $\hat{\mu}_{k}=\frac{1}{N_{k}}\sum_{Z_{i}(\tau_{u})=k}Y_{i}$ and $\hat{p}_{k}=\frac{N_{k}}{N_{0}+N_{1}}$ be the sample means and treatment shares. Define the random interval
$\begin{split}\mathcal{H}_{\alpha_{u}}[\Re_{u}]=\left[\hat{\mu}_{1}^{-}\hat{p}_{1}^{-}+L_{\alpha_{u}}^{(1)}\hat{p}_{0}^{-}-U_{\alpha_{u}}^{(0)}\hat{p}_{1}^{+}-\hat{\mu}_{0}^{+}\hat{p}_{0}^{+},\;\hat{\mu}_{1}^{+}\hat{p}_{1}^{+}+U_{\alpha_{u}}^{(1)}\hat{p}_{0}^{+}-L_{\alpha_{u}}^{(0)}\hat{p}_{1}^{-}-\hat{\mu}_{0}^{-}\hat{p}_{0}^{-}\right],\end{split}$ (30)
where $\hat{\mu}_{k}^{\pm}=\hat{\mu}_{k}\pm t_{\mu,k}$ and $\hat{p}_{k}^{\pm}=\hat{p}_{k}\pm t_{p,k}$. Then, under assumption 1
$\Pr\left(\forall u\in\mathcal{M},\;\Re_{u}\in\mathcal{H}_{\alpha_{u}}[\Re_{u}]\right)\geq 1-\alpha,$ (31)
Proposition 2 states that the data-driven set $\mathcal{H}_{\alpha_{u}}[\Re_{u}]$ in (30) is a $100(1-\alpha_{u})\%$-level confidence region for the average treatment effect $\Re_{u}$ under nothing more than i.i.d. sampling. Because $\mathbb{E}[Y_{ijt}^{(1)}\mid Z_{ijt}(\tau_{u})=0]$ and $\mathbb{E}[Y_{ijt}^{(0)}\mid Z_{ijt}(\tau_{u})=1]$ are latent, point identification is impossible without additional assumptions; the proposition nevertheless guarantees that the random interval constructed from the empirical means, treatment proportions, and DKW-calibrated outward empirical quantiles will cover the true $\Re_{u}$ in at least $100(1-\alpha_{u})\%$ of repeated samples. Practically, one computes $\mathcal{H}_{\alpha_{u}}[\Re_{u}]$ by (i) splitting the sample into treated and untreated subsamples, (ii) forming the subsample-specific means $\hat{\mu}_{k}$ and proportions $\hat{p}_{k}$, (iii) computing $\varepsilon_{k}=\sqrt{\log(12/\alpha_{u})/(2N_{k})}$ and choosing any $r_{k}\in(0,\frac{1}{2}-\varepsilon_{k}]$, then setting the tail endpoints
$L_{\alpha_{u}}^{(k)}=Y^{(k)}_{\left(\lceil(r_{k}+\varepsilon_{k})N_{k}\rceil\right)},\quad U_{\alpha_{u}}^{(k)}=Y^{(k)}_{\left(\lceil(1-r_{k}-\varepsilon_{k})N_{k}\rceil\right)},$
(in particular, if $\varepsilon_{k}\leq\frac{1}{4}$ one may take $r_{k}=\varepsilon_{k}$, yielding the convenient indices $\lceil 2\varepsilon_{k}N_{k}\rceil$ and $\lceil(1-2\varepsilon_{k})N_{k}\rceil$), and (iv) plugging these objects into (30). The resulting band can be used exactly like an ordinary confidence interval: the null hypothesis $H_{0}:0\in\Re_{u}$ is rejected at level $\alpha_{u}$ whenever $0\notin\mathcal{H}_{\alpha_{u}}[\Re_{u}]$.
If substantive knowledge implies that the latent outcomes are truncated on one or both tails, the extreme-value inputs in Manski’s bounds can be replaced by the true population limits. Let $\lambda$ (lower) and $\Lambda$ (upper) denote any such known bounds. When both limits are known one sets $L^{(k)}=\lambda$ and $U^{(k)}=\Lambda$ in the plug-in formulas; the resulting $100(1-\alpha)\%$ simultaneous band coincides with Proposition 1 and requires no DKW calibration. When only one tail is known, say $Y\geq\lambda$, but the upper support is unknown, we fix the lower extreme at $\lambda$ while retaining the DKW-calibrated endpoint on the upper side. The next Corollary shows that this hybrid construction preserves the nominal family-wise coverage probability even when the first significant threshold is data-selected.
Corollary 1 (Finite-sample coverage under i.i.d. sampling and truncated distribution).
Let $0<\alpha_{u}<1$ for $u=m_{0},\cdots,m_{1}$ and denote by $N_{k}=\sum_{i}\mathbbm{1}\{Z_{i}(\tau_{u})=k\}$ the sample size in treatment group $k\in\{0,1\}$ and by
$\lambda\leq Y^{(k)}_{(1)}\leq\cdots\leq Y^{(k)}_{(N_{k})}$
the order statistics of the observed outcomes in that group. Set
$\varepsilon_{k}:=\sqrt{\frac{\log\left(6/\alpha_{u}\right)}{2N_{k}}},\quad L^{(k)}:=\lambda,$
choose any $r_{k}\in(0,\,1-\varepsilon_{k}]$, and define the upper endpoint via the one-sided DKW quantile relation as
$U_{\alpha_{u}}^{(k)}:=\hat{F}^{-1}_{N_{k}}(1-r_{k}-\varepsilon_{k})=Y^{(k)}_{\left(\lceil(1-r_{k}-\varepsilon_{k})N_{k}\rceil\right)},\qquad k=0,1.$
Define the two thresholds
@@
$\begin{split}\mathcal{H}_{\alpha_{u}}[\Re_{u}]=\left[\hat{\mu}_{1}^{-}\hat{p}_{1}^{-}+L^{(1)}\hat{p}_{0}^{-}-U_{\alpha_{u}}^{(0)}\hat{p}_{1}^{+}-\hat{\mu}_{0}^{+}\hat{p}_{0}^{+},\;\hat{\mu}_{1}^{+}\hat{p}_{1}^{+}+U_{\alpha_{u}}^{(1)}\hat{p}_{0}^{+}-L^{(0)}\hat{p}_{1}^{-}-\hat{\mu}_{0}^{-}\hat{p}_{0}^{-}\right],\end{split}$ (32)
where $\hat{\mu}_{k}^{\pm}=\hat{\mu}_{k}\pm t_{\mu,k}$ and $\hat{p}_{k}^{\pm}=\hat{p}_{k}\pm t_{p,k}$. Then, under assumption 1
$\Pr\left(\forall u\in\mathcal{M},\;\Re_{u}\in\mathcal{H}_{\alpha_{u}}[\Re_{u}]\right)\geq 1-\alpha,$ (33)
@@
Proposition 3 (Finite-sample coverage under $\upalpha$-mixing sampling).
Let $0<\alpha_{u}<1$ for $u=m_{0},\cdots,m_{1}$ and let $(Y_{ijt},Z_{jt})_{i,j,t}$ be a strictly stationary sequence with $Z_{i}(\tau_{u})\in\{0,1\}$, $Y_{i}\in\mathbb{R}$, and strong-mixing coefficients $\upalpha(r)$ satisfying
@@
$N_{k}=\sum_{i=1}^{n}\mathbbm{1}\{Z_{i}(\tau_{u})=k\},\quad\hat{p}_{k}=\frac{N_{k}}{N_{0}+N_{1}},\quad\hat{\mu}_{k}=\frac{1}{N_{k}}\sum_{Z_{i}(\tau_{u})=k}Y_{i},\quad k=0,1.$
@@
$t_{p,k}:=(1+4C_{\upalpha})\sqrt{\frac{2\log(12/\alpha_{u})}{N_{k}}},\qquad t_{\mu,k}:=\max\left\{t_{k}^{(1)},t_{k}^{(2)},t_{k}^{(3)}\right\},$
@@
$\varepsilon_{k}:=t_{p,k},$
choose any $r_{k}\in\left(0,\frac{1}{2}-\varepsilon_{k}\right]$, and define the DKW-calibrated endpoints
$L^{(k)}_{\alpha_{u}}:=Y^{(k)}_{\left(\lceil(r_{k}+\varepsilon_{k})N_{k}\rceil\right)},\quad U^{(k)}_{\alpha_{u}}:=Y^{(k)}_{\left(\lceil(1-r_{k}-\varepsilon_{k})N_{k}\rceil\right)},\quad k=0,1.$
(In particular, if $\varepsilon_{k}\leq\frac{1}{4}$, one may take $r_{k}=\varepsilon_{k}$, yielding $L^{(k)}_{\alpha_{u}}=Y^{(k)}_{(\lceil 2\varepsilon_{k}N_{k}\rceil)}$ and $U^{(k)}_{\alpha_{u}}=Y^{(k)}_{(\lceil(1-2\varepsilon_{k})N_{k}\rceil)}$.)
@@
$\mathcal{H}_{\alpha_{u}}[\Re_{u}]=\left[\hat{\mu}_{1}^{-}\hat{p}_{1}^{-}+L^{(1)}_{\alpha_{u}}\hat{p}_{0}^{-}-U^{(0)}_{\alpha_{u}}\hat{p}_{1}^{+}-\hat{\mu}_{0}^{+}\hat{p}_{0}^{+},\;\hat{\mu}_{1}^{+}\hat{p}_{1}^{+}+U^{(1)}_{\alpha_{u}}\hat{p}_{0}^{+}-L^{(0)}_{\alpha_{u}}\hat{p}_{1}^{-}-\hat{\mu}_{0}^{-}\hat{p}_{0}^{-}\right],$
where $\hat{\mu}_{k}^{\pm}=\hat{\mu}_{k}\pm t_{\mu,k}$ and $\hat{p}_{k}^{\pm}=\hat{p}_{k}\pm t_{p,k}$. Then, under assumption 2
$\Pr\left(\forall u\in\mathcal{M},\;\Re_{u}\in\mathcal{H}_{\alpha_{u}}[\Re_{u}]\right)\geq 1-\alpha.$
Similar to Corollary 1, the extension of Proposition 3 to the case of one-tail truncated latent conditional expectations simply requires modifying the endpoints. Specifically, for the case of lower-tail truncation, replace
$L^{(k)}:=\lambda,\quad U^{(k)}_{\alpha_{u}}:=\hat{F}^{-1}_{N_{k}}\left(1-r_{k}-\tilde{\varepsilon}_{k}\right)=Y^{(k)}_{\left(\lceil(1-r_{k}-\tilde{\varepsilon}_{k})\,N_{k}\rceil\right)},$
with the fixed lower bound and a one-sided DKW-calibrated upper empirical quantile, where
$\tilde{\varepsilon}_{k}:=(1+4C_{\upalpha})\sqrt{\frac{2\log\!\left(6/\alpha_{u}\right)}{N_{k}}},\qquad r_{k}\in\left(0,1-\tilde{\varepsilon}_{k}\right],$
for $k=0,1$, and the observed order statistics satisfy $\lambda\leq Y_{(1)}^{(k)}\leq\cdots\leq Y_{\left(\lceil(1-r_{k}-\tilde{\varepsilon}_{k})(N_{k})\rceil\right)}^{(k)}$. (When $\tilde{\varepsilon}_{k}\leq\tfrac{1}{4}$, a convenient choice is $r_{k}=\tilde{\varepsilon}_{k}$, yielding the index $\lceil(1-2\tilde{\varepsilon}_{k})N_{k}\rceil$.)